OpenStudy (elifbestegul):

4sinx - 3cosx = 0 is given. Find sin2x.

OpenStudy (kashmoneyjr):

Too lazy to solve it by myself lol, but heres my take: Add 3cosx to each side 4sinx=3cosx divide each side by cosx to get 4(sinx/cosx)=3 divide each side by 4 to get sinx/cosx=3/4 sinx/cosx=tanx, so tanx=3/4 x=inversetan(3/4) (which im too lazy to calculate) then calculate sin2x, using the value you found

OpenStudy (elifbestegul):

Thank you but I was not looking for an answer which would need to be calculated on a calculator. Thanks anyway @kashmoneyjr

OpenStudy (kashmoneyjr):

Thats a tough question then you probably already know this, but using the identity sin2x=2sinxcosx may help

OpenStudy (elifbestegul):

yes, it needs to be done with that formula but I can't figure out how to @kashmoneyjr

OpenStudy (irishboy123):

\(4 \sin x = 3 \cos x \)

OpenStudy (irishboy123):

\(\tan x = \dfrac{3}{4}\) 

OpenStudy (kashmoneyjr):

^this: use the 3-4-5 triangle

OpenStudy (kashmoneyjr):

if tanx=3/4, then by the right triangle definition, sinx=3/5 and cosx=4/5 so sin2x=2sinxcosx=2*(3/5)*(4/5)

OpenStudy (kashmoneyjr):

so 24/25

OpenStudy (irishboy123):